# GNSS-based Lunar Orbit and Clock Estimation With Stochastic Cloning UD Filter

### v2

@@ 1 Introduction
Third, the effects of general and special relativity on GNSS signals and the local clocks of lunar satellites have not been adequately addressed in prior GNSS-based lunar orbit determination and clock synchronization (ODTS) works. Relativistic effects induce coupling between orbital states and clock bias states, leading to systematic biases in measurements and estimated clock biases if not properly modeled [seyffert2025]. While the LuGRE mission synchronized its clock to GPS Time (GPST) [Konitzer2022LuGRE], future lunar satellites aiming for high-precision PNT services will need to synchronize clocks to lunar-based reference time frames [Bourgoin_2025, Turyshev2026RelativisticTimescales, Kopeikin_2024], making accurate modeling of these relativistic effects essential.
@@ 2.1 State Definition
$X_{k}=\begin{bmatrix}\mathbf{r}_{k}^{\mathrm{LCRF}}&\mathbf{v}_{k}^{\mathrm{LCRF}}&c\,\delta t_{k}&c\,\delta\dot{t}_{k}&c\,\delta\ddot{t}_{k}&\gamma_{\mathrm{srp}}\end{bmatrix}^{\!T},$ (2)
@@ 2.2 Orbit Dynamics
$\mathcal{X}^{\mathrm{TCL}}$ $=\mathcal{X}^{\mathrm{TCB}}=\frac{1}{1-L_{B}}\,\mathbf{X}^{\mathrm{TDB}},$ (6)
@@ 2.3.1 Lunar Coordinate Time (TCL) and Lunar Time (LT)
TCL is the relativistic coordinate time associated with the LCRS, analogous to Geocentric Coordinate Time (TCG) for the GCRS and Barycentric Coordinate Time (TCB) for the BCRS. The transformation between TCL and TCB is defined by accounting for the Moon’s barycentric motion and the external gravitational potential acting on the Moon:
$\frac{d\mathrm{TCL}}{d\mathrm{TCB}}=1-\frac{1}{c^{2}}\left(\frac{\|\mathbf{v}_{m}\|^{2}}{2}+U_{\mathrm{ext}}(\mathbf{x}_{m})\right)+\mathcal{O}\!\left(\frac{1}{c^{4}}\right),$ (8)
where $\mathbf{x}_{m}$ and $\mathbf{v}_{m}$ denote the Moon’s barycentric position and velocity, respectively, and $U_{\mathrm{ext}}(\mathbf{x}_{m})$ represents the gravitational potential of all bodies excluding the Moon, evaluated at the Moon’s location.
@@
where $L_{G}=6.969290134\times 10^{-10}$ and $L_{L}$ are linear scale constants relating TCG to TT and TCL to LT, respectively, and $T_{0}=2443144.5003725$ JD (Julian Date) is the reference epoch. The term $\Delta(\mathrm{TCG}-\mathrm{TCL})$ is given by (9). The coefficient $L_{L}=\Phi_{L}/c^{2}$ is depends on the reference chosen lunar potential $\Phi_{L}$. Reference [Kopeikin_2024] adopts a lunar gravitational potential of $\Phi_{L}=2.822336927\times 10^{6},\mathrm{m}^{2}/\mathrm{s}^{2}$, yielding $L_{L}=3.14027\times 10^{-11}$, whereas [Turyshev2026RelativisticTimescales] assumes $\Phi_{L}=2.82123744381\times 10^{6},\mathrm{m}^{2}/\mathrm{s}^{2}$, resulting in $L_{L}=3.13905\times 10^{-11}$. The definitive value of $L_{L}$ has not yet been established through international standardization, and the initial lunar timescale offset, $T_{L0}$, likewise remains undefined.
@@ 2.3.2 Relativistic Clock Drift with respect to TCL
$\dot{\delta}t_{\mathrm{rel}}=\frac{d(\tau-\mathrm{TCL})}{d\mathrm{TCL}}=\frac{1}{c^{2}}\left(U_{L}(r)+\frac{v^{2}}{2}\right)\approx\frac{1}{c^{2}}\left(\frac{\mu_{L}}{r}+\frac{v^{2}}{2}\right),$ (11)
where $\tau$ denotes the satellite proper time, $U_{L}(r)$ is the lunar gravitational potential at the spacecraft location, $r$ is the geocentric distance, and $v$ is the orbital velocity in the LCRS.
@@ 2.3.4 Stochastic Clock Drift Model
$\begin{bmatrix}c\delta t_{k}\\ c\delta\dot{t}_{k}\\ c\delta\ddot{t}_{k}\end{bmatrix}$ $=\begin{bmatrix}1&\Delta t&\frac{1}{2}\Delta t^{2}\\ 0&1&\Delta t\\ 0&0&1\end{bmatrix}\begin{bmatrix}c\delta t_{k-1}\\ c\delta\dot{t}_{k-1}\\ c\delta\ddot{t}_{k-1}\end{bmatrix}+\begin{bmatrix}c\int_{t_{k-1}}^{t_{k}}\dot{\delta}t_{\mathrm{rel}}\,dt\\ c\dot{\delta}t_{\mathrm{rel}}\\ 0\end{bmatrix}+\boldsymbol{\epsilon}_{k}^{\mathrm{clk}},$ (13)
where $\Delta t=t_{k}-t_{k-1}$ and $\boldsymbol{\epsilon}_{k}^{\mathrm{clk}}\sim\mathcal{N}(0,Q_{k}^{\mathrm{clk}})$ represents the clock process noise. This formulation explicitly couples clock and orbital dynamics through the relativistic drift term $\dot{\delta}t_{\mathrm{rel}}$.
@@ 4.1.1 Time Update
$f^{\text{dyn}}(t,\mathbf{x})=\frac{d}{dt}\mathbf{x}$ $=\frac{d}{dt}\begin{bmatrix}\mathbf{r}_{k}&\mathbf{v}_{k}&c\delta t_{k}&c\delta\dot{t}_{k}&c\delta\ddot{t}_{k}&\gamma_{k}\end{bmatrix}^{\top}$ (58)
@@ 4.1.2 Measurement Preprocessing
    $\begin{cases}\text{Accept TDCP measurement if}&|z_{k,i}-\hat{z}_{k,i}|<3\sqrt{\mathbf{H}_{k,i}\hat{P}_{i,i}\mathbf{H}_{k,i}^{T}+R_{k,i}}<\lambda_{L}\\ \text{Discard TDCP measurement if}&\text{otherwise}\end{cases}$ (66)
    This criterion discards the TDCP measurements whenever the residual exceeds a 3-sigma bound, or exceeds the wavelength of the signal $\lambda_{L}$ (=19 cm for L1 and 24 cm for L5). It also discards all measurements where the 3-sigma bound itself exceeds the wavelength, which indicates insufficient state knowledge and/or measurement noise to reliably detect cycle slips. For this work, we only discard the TDCP measurements with detected cycle slips, and do not attempt to repair them.
@@ 5.2 GNSS Orbits and Ephemerides
We considered a multi-constellation set of GPS, Galileo, and Quasi-Zenith Satellite System (QZSS) satellites. Truth GNSS orbits were generated using the International GNSS Service (IGS) final orbit products [Johnston2017]. Specifically, final orbit products provide satellite center-of-mass positions, which were interpolated using cubic splines to obtain satellite positions at the desired epochs. Antenna phase center information from ANTEX files was then used to map center-of-mass positions to antenna reference point (ARP) positions.
To emulate real-time navigation message performance, broadcast ephemerides were retrieved from daily multi-GNSS broadcast ephemeris files via Crustal Dynamics Data Information System (CDDIS). The broadcast ephemerides were used within the measurement models (Section 2) to introduce realistic ephemeris and clock modeling errors relative to the truth orbits. The resulting GNSS constellation geometry at the simulation epoch is shown in Figure 5.
@@ 6.2 Satellite Ephemeris Errors
The ephemeris error distributions for the GPS (L1/L5), Galileo (E1/E5a), and QZSS (L1/L5) satellites in directions along the GNSS-to-Lunar satellite line-of-sight is shown in Fig. 10. We see that most of the position and clock bias ephemeris errors are within $\pm 2$ m. The total user range error from the ephemeris are mostly within $\pm 4$ m, and has different mean biases for different GNSS constellations. The mean and variances of the ephemeris errors in LOS direction for different GNSS constellations are summarized in Table 6.
@@ 6.3 Receiver Noise
Figure 11: Lunar GNSS receiver receiver noise vs tangential altitude
@@ 6.4 Ionospheric and Plasmaspheric Delays
Figure 12: Ionospheric and plasmaspheric delays vs tangential altitude
@@ 7.1 Filter Settings
Based on the measurement error analysis in Section 6, we set the broadcast-ephemeris-induced range error standard deviation to $\sigma_{\text{URE}}=$5.0\text{\,}\mathrm{m}$\$ for pseudorange and the time-differenced equivalent to $\sigma_{\Delta\text{URE}}=$8\text{\,}\mathrm{m}\mathrm{m}$\$ for TDCP. The altitude masks and these inflation terms were selected from multiple trial runs; further tuning may yield additional improvement.
@@
For each measurement configuration, we run 50 Monte-Carlo simulations with independent realizations of the initial state perturbations and measurement errors. To ensure convergence from the large initial uncertainties, performance is evaluated over the final orbit revolution (approximately 30 hours). We report SISE metrics consistent with the LunaNet SRD [nasa2022srd] as follows:
@@ 7.2 Results
Table 9 summarizes the position and velocity SISE statistics (RMS, 95%, and 99.7% percentiles) for the filter and smoother, averaged over 50 Monte-Carlo runs.
Table 9: SISE position and velocity errors of the filter and smoother at the final orbit under different measurement configurations. Values are the means of over 50 Monte-Carlo simulations. (PR=pseudorange, TDCP=time-differenced carrier phase, IF=ionosphere-free).
@@
Filter 21.41 32.60 36.17 1.16 2.02 2.66
($\geq$ 1000 km) Smoother 12.97 16.11 16.18 0.94 1.96 2.26
Filter 4.38 6.82 8.11 0.43 0.86 1.26
($\geq$ 1000 km) Smoother 4.55 6.21 6.89 0.35 0.80 1.10
Filter 3.68 5.79 6.79 0.39 0.77 1.90
($\geq$ 1000 km) ($\geq$ 5000 km) Smoother 3.89 5.31 5.68 0.25 0.54 0.80
@@ 7.2.3 Ionosphere-free Pseudorange + TDCP
Adding TDCP further improves SISE performance for both the filter and smoother. TDCP provides high-precision relative information between successive epochs, tightening constraints on velocity- and clock-drift-related states and reducing overall estimation error. Relative to ionosphere-free pseudorange alone, the filter RMS SISE improves by approximately 16–18% for position and 9–11% for velocity.
@@ 7.2.4 Effect of Smoothing
Applying the fixed-interval smoother improves both position and velocity SISE for all measurement configurations. The TDCP-assisted smoother achieves the best overall performance. Figure 15 compares the filter and smoother trajectories (after five filter–smoother iterations) for one Monte-Carlo run with ionosphere-free pseudorange and TDCP. Large position SISE errors near apolune, where slower orbital motion reduces instantaneous observability, and larger velocity SISE errors near perilune are both noticeably reduced after smoothing.
Smoothing also suppresses the occasional filter spikes induced by near-outlier TDCP updates, reducing the 99.7% velocity SISE from $1.90\text{\,}\mathrm{m}\mathrm{m}\mathrm{/}\mathrm{s}$ to $0.80\text{\,}\mathrm{m}\mathrm{m}\mathrm{/}\mathrm{s}$ (58% reduction). Overall, the smoothed velocity SISE with TDCP is approximately 30% lower than the ionosphere-free pseudorange-only smoother.

### v3

@@ 1 Introduction
Third, the effects of general and special relativity on GNSS signals and the local clocks of lunar satellites have not been adequately addressed in prior GNSS-based lunar orbit determination and clock synchronization (ODTS) works. Relativistic effects induce coupling between orbital states and clock bias states, leading to systematic drift (in orders of $mu$ seconds per day) in estimated clock biases if not properly modeled [seyffert2025]. In addition, while the LuGRE mission synchronized its clock to GPS Time (GPST) [Konitzer2022LuGRE], future lunar satellites aiming for high-precision PNT services will need to synchronize clocks to lunar-based reference time frames [Bourgoin_2025, Turyshev2026RelativisticTimescales, Kopeikin_2024], making accurate conversions between these time scales essential.
@@ 2.1 State Definition
$\mathbf{x}_{k}=\begin{bmatrix}\mathbf{r}_{k}^{\mathrm{LCRF}}&\mathbf{v}_{k}^{\mathrm{LCRF}}&c\,\delta t_{k}&c\,\delta\dot{t}_{k}&c\,\delta\ddot{t}_{k}&\gamma_{\mathrm{srp}}\end{bmatrix}^{\!T},$ (2)
@@ 2.2 Orbit Dynamics
$\mathcal{X}^{\mathrm{TCL}}$ $=\mathcal{X}^{\mathrm{TCB}}=\frac{1}{1-L_{B}}\,\mathcal{X}^{\mathrm{TDB}},$ (6)
@@ 2.3.1 Lunar Coordinate Time (TCL) and Lunar Time (LT)
TCL is the relativistic coordinate time associated with the LCRS, analogous to Geocentric Coordinate Time (TCG) for the Geocentric Celestial Reference System (GCRS) and Barycentric Coordinate Time (TCB) for the Barycentric Celestial Reference System (BCRS). The transformation between TCL and TCB is defined by accounting for the Moon’s barycentric motion and the external gravitational potential acting on the Moon:
$\frac{d\mathrm{TCL}}{d\mathrm{TCB}}=1-\frac{1}{c^{2}}\left(\frac{\|\mathbf{v}_{m}\|^{2}}{2}+U_{\mathrm{ext}}(\mathbf{r}_{m})\right)+\mathcal{O}\!\left(\frac{1}{c^{4}}\right),$ (8)
where $\mathbf{r}_{m}$ and $\mathbf{v}_{m}$ denote the Moon’s barycentric position and velocity, respectively, and $U_{\mathrm{ext}}(\mathbf{x}_{m})$ represents the gravitational potential of all bodies excluding the Moon, evaluated at the Moon’s location.
@@
where $L_{G}=6.969290134\times 10^{-10}$ and $L_{L}$ are linear scale constants relating TCG to TT and TCL to LT, respectively, and $T_{0}=2443144.5003725$ JD (Julian Date) = 1977-01-01 00:00:32.184 TAI is the reference epoch. The term $\Delta(\mathrm{TCL}-\mathrm{TCG})$ is given by (9). The coefficient $L_{L}=\Phi_{L}/c^{2}$ is depends on the reference chosen lunar potential $\Phi_{L}$. Reference [Kopeikin_2024] adopts a lunar gravitational potential of $\Phi_{L}=2.822336927\times 10^{6},\mathrm{m}^{2}/\mathrm{s}^{2}$, yielding $L_{L}=3.14027\times 10^{-11}$, whereas [Turyshev2026RelativisticTimescales] assumes $\Phi_{L}=2.82123744381\times 10^{6},\mathrm{m}^{2}/\mathrm{s}^{2}$, resulting in $L_{L}=3.13905\times 10^{-11}$. The definitive value of $L_{L}$ has not yet been established through international standardization, and the initial lunar timescale offset, $T_{L0}$, likewise remains undefined.
@@ 2.3.2 Relativistic Clock Drift with respect to TCL
$\dot{\delta}t_{\mathrm{rel}}(\mathbf{r},\mathbf{v})=\frac{d(\tau-\mathrm{TCL})}{d\mathrm{TCL}}=\frac{1}{c^{2}}\left(U_{L}(\mathbf{r})+\frac{v^{2}}{2}\right)\approx\frac{1}{c^{2}}\left(\frac{\mu_{L}}{r}+\frac{v^{2}}{2}\right),$ (11)
where $\tau$ denotes the satellite proper time, $U_{L}(\mathbf{r})$ is the lunar gravitational potential at the spacecraft location $\mathbf{r}$, $r$ is the lunicentric distance, and $\mathbf{v},v$ is the orbital velocity and its norm in the LCRS.
@@ 2.3.4 Stochastic Clock Drift Model
$\begin{bmatrix}c\delta t_{k}\\ c\delta\dot{t}_{k}\\ c\delta\ddot{t}_{k}\end{bmatrix}$ $=\begin{bmatrix}1&\Delta t&\frac{1}{2}\Delta t^{2}\\ 0&1&\Delta t\\ 0&0&1\end{bmatrix}\begin{bmatrix}c\delta t_{k-1}\\ c\delta\dot{t}_{k-1}\\ c\delta\ddot{t}_{k-1}\end{bmatrix}+\begin{bmatrix}c\int_{t_{k-1}}^{t_{k}}\dot{\delta}t_{\mathrm{rel}}(\mathbf{r}(t^{\prime}),\mathbf{v}(t^{\prime}))dt^{\prime}\\ c\dot{\delta}t_{\mathrm{rel}}(\mathbf{r}(t_{k}),\mathbf{v}(t_{k}))\\ 0\end{bmatrix}+\boldsymbol{\epsilon}_{k}^{\mathrm{clk}},$ (13)
where $\Delta t=t_{k}-t_{k-1}$ and $\boldsymbol{\epsilon}_{k}^{\mathrm{clk}}\sim\mathcal{N}(0,Q_{k}^{\mathrm{clk}})$ represents the clock process noise. This formulation explicitly couples clock and orbital dynamics through the relativistic drift term $\dot{\delta}t_{\mathrm{rel}}((\mathbf{r}(t_{k}),\mathbf{v}(t_{k}))$.
@@ 4.1.1 Time Update
$f^{\text{dyn}}(t,\mathbf{x})$ $=\frac{d}{dt}\mathbf{x}$ (58)
$=\frac{d}{dt}\begin{bmatrix}\mathbf{r}_{k}&\mathbf{v}_{k}&c\delta t_{k}&c\delta\dot{t}_{k}&c\delta\ddot{t}_{k}&\gamma_{k}\end{bmatrix}^{\top}$ (59)
@@ 4.1.2 Measurement Preprocessing
    $\begin{cases}\text{Accept TDCP measurement}\ i\ \text{if}&|z_{k,i}-\hat{z}_{k,i}|<2.5S_{k,i},\ \text{and}\ 3S_{k,i}<\lambda_{L}\\ \text{Discard TDCP measurement}\ i\ \text{if}&\text{otherwise}\end{cases}$ (67)
    where $S_{k,i}=\sqrt{\mathbf{H}_{k,i}\hat{P}_{i,i}\mathbf{H}_{k,i}^{T}+R_{k,i}}$. This criterion discards the TDCP measurements whenever the residual exceeds a 2.5-sigma bound, or exceeds the wavelength of the signal $\lambda_{L}$ (=19 cm for L1 and 24 cm for L5). It also discards all measurements where the 3-sigma bound itself exceeds the wavelength, which indicates insufficient state knowledge and/or measurement noise to reliably detect cycle slips. For this work, we only discard the TDCP measurements with detected cycle slips, and do not attempt to repair them.
@@ 5.2 GNSS Orbits and Ephemerides
We considered a multi-constellation set of GPS, Galileo, and Quasi-Zenith Satellite System (QZSS) satellites. Truth GNSS orbits were generated using the International GNSS Service (IGS) final orbit products [Johnston2017]. Specifically, final orbit products provide satellite center-of-mass positions, which were interpolated using cubic splines to obtain satellite positions at the desired epochs. Antenna phase center information from ANTEX files was then used to map center-of-mass positions to antenna reference point (ARP) positions. The resulting GNSS constellation geometry at the simulation epoch is shown in Figure 5.
To emulate real-time navigation message performance, broadcast ephemerides were retrieved from daily multi-GNSS broadcast ephemeris files via Crustal Dynamics Data Information System (CDDIS). The broadcast ephemerides were used within the measurement models (Section 2) to introduce realistic ephemeris and clock modeling errors relative to the truth orbits.
In addition, to remove the systematic bias common to all satellite clocks arising from differences in the time references used by the broadcast and precise clock products, the precise clock bias used as truth is adjusted as [Wang_2019]
$\delta t^{\mathrm{pr}}_{k}\leftarrow\delta t^{\mathrm{pr}}_{k}-\operatorname{median}\left(\delta t^{\mathrm{pr}}-\delta t^{\mathrm{br}}\right),$ (73)
where $\delta t^{\mathrm{br}}_{k}$ is the broadcast clock bias at time step $k$, and the median is computed over the entire simulation interval for each constellation. For QZSS satellites, an analogous correction is applied to orbits on a per-satellite basis to remove the constant offset between the precise and broadcast ephemerides in the radial direction.
@@ 6.2 Satellite Ephemeris Errors
The ephemeris error distributions for the GPS (L1/L5), Galileo (E1/E5a), and QZSS (L1/L5) satellites in directions along the GNSS-to-Lunar satellite line-of-sight is shown in Fig. 10. We see that most of the position and clock bias ephemeris errors are within $\pm 2$ m. The total user range error from the ephemeris are mostly within $\pm 3$ m, and has different mean biases for different GNSS constellations. The mean and variances of the ephemeris errors in LOS direction for different GNSS constellations are summarized in Table 6.
@@ 6.3 Receiver Noise
Figure 11: GNSS receiver noise vs tangential altitude for the L1 pseudorange (left), TDCP (center), and ionosphere-free combination (right), evaluated over a 180-hour simulation for an LCRNS satellite. The main lobe signals ($\leq$ 2000 km) have lower noise compared to the sidelobe signals with lower $C/N_{0}$.
@@ 6.4 Ionospheric and Plasmaspheric Delays
Figure 12: Ionospheric and plasmaspheric group delays vs tangential altitude for the L1 pseudorange (left), TDCP (center), and ionosphere-free combination (right), evaluated over a 180-hour simulation for an LCRNS satellite. For signal paths intersecting the ionosphere (tangential altitude $\leq 1000$ km), higher-order ionospheric effects and ray-bending delays, which are not fully eliminated by the ionosphere-free combination, remain observable (right).
@@ 7.1 Filter Settings
Based on the measurement error analysis in Section 6, we set the broadcast-ephemeris-induced range error standard deviation to $\sigma_{\text{URE}}=$10.0\text{\,}\mathrm{m}$\$ for pseudorange and the time-differenced equivalent to $\sigma_{\Delta\text{URE}}=$8\text{\,}\mathrm{m}\mathrm{m}$\$ for TDCP. The altitude masks and these inflation terms were selected from multiple trial runs; further tuning may yield additional improvement.
@@
For each measurement configuration, we run 40 Monte-Carlo simulations with independent realizations of the initial state perturbations and measurement errors. To ensure convergence from the large initial uncertainties, performance is evaluated over the final orbit revolution (approximately 30 hours). We report SISE metrics consistent with the LunaNet SRD [nasa2022srd] as follows:
@@ 7.2 Results
Table 9 summarizes the position and velocity SISE statistics (RMS, 95%, and 99.7% percentiles) for the filter and smoother, averaged over 40 Monte-Carlo runs.
Table 9: SISE position and velocity errors of the filter and smoother at the final orbit under different measurement configurations. Values represent are computed over 40 Monte Carlo simulations (PR: pseudorange, TDCP: time-differenced carrier phase, IF: ionosphere-free). Using L1 pseudorange only results in the largest SISE errors due to unmodeled plasmaspheric delays. The combination of ionosphere-free pseudorange and TDCP achieves the smallest SISE errors. Application of smoothing reduces the 95th- and 99.7th-percentile SISE errors.
@@
Filter 18.41 28.73 31.31 0.82 1.48 2.11
($\geq$ 1000 km) Smoother 10.84 14.26 14.85 0.70 1.58 2.13
Filter 6.05 9.74 12.39 0.46 0.91 1.80
($\geq$ 1000 km) Smoother 4.77 7.71 9.65 0.36 0.74 1.54
Filter 4.53 7.44 10.20 0.45 0.98 2.08
($\geq$ 1000 km) ($\geq$ 5000 km) Smoother 2.79 4.47 7.31 0.30 0.64 1.46
@@ 7.2.3 Ionosphere-free Pseudorange + TDCP
Adding TDCP to the ionosphere-free pseudorange improves the overall SISE performance for both the filter and smoother. For the filter, the position RMS decreases from 6.05m (IF PR only) to 4.53m (IF PR + TDCP), corresponding to an improvement of approximately 25%. The position 95% and 99.7% SISE are also reduced, from 9.74m and 12.39m to 7.44m and 10.20m, respectively. These results confirm that TDCP provides strong relative position and clock bias constraints between successive epochs, particularly benefiting the position solution.
@@ 7.2.4 Effect of Smoothing
Applying the fixed-interval smoother improves both position and velocity SISE across all measurement configurations. The largest benefit is observed when TDCP is included. With ionosphere-free pseudorange only, smoothing reduces the position RMS from 6.05m to 4.77m and the velocity RMS from 0.46mm/s to 0.36mm/s. When TDCP is added, the improvements are more pronounced: the position RMS decreases from 4.53m to 2.79m (approximately 38% reduction), and the velocity RMS decreases from 0.45mm/s to 0.30mm/s (approximately 33% reduction).
Figure 15 compares the filter and smoother trajectories (after five filter–smoother iterations) for one Monte-Carlo run with ionosphere-free pseudorange and TDCP. Large position SISE errors near apolune, where slower orbital motion reduces instantaneous observability, and larger velocity SISE errors near perilune are both noticeably reduced after smoothing.
Smoothing also mitigates the high-percentile degradation introduced by TDCP in the filter. In particular, the 99.7% velocity SISE decreases from 2.08mm/s (filter) to 1.46mm/s (smoother), a reduction of about 30%. Compared to the ionosphere-free pseudorange-only smoother (1.54mm/s at 99.7%), the TDCP-assisted smoother achieves slightly lower extreme velocity errors, while providing substantially improved position accuracy (2.79m vs. 4.77 m RMS). These results indicate that the backward information flow in the smoother effectively suppresses transient TDCP-induced spikes and yields the best overall consistency and accuracy among the tested configurations.